OpenStudy (anonymous):

Find the area of the region. Use a graphing utility to verify your result. y = 6 sin(x) + sin(6x) the graph goes from x=0 to x=pi

OpenStudy (anonymous):

can someone please helpme

OpenStudy (anonymous):

satelite can u help me i knwo i ahve to find the integral first but then what

OpenStudy (anonymous):

...

OpenStudy (anonymous):

\[\int_0^{\pi} 6\sin(x)+\sin(6x)dx\]

OpenStudy (anonymous):

you want a graphing utility? here: http://www.wolframalpha.com/input/?i=integra+0+to+pi+%286sin%28x%29%2Bsin%286x%29%29+dx

OpenStudy (anonymous):

Help me satellite73

OpenStudy (anonymous):

answer seems to be 12 but the integral is easy enough. antiderivative of \[6\sin(x)\] is \[-6\cos(x)\] and antiderivative of \[\sin(6x)\] is \[-\frac{1}{6}\cos(6x)\] so \[\int 6\sin(x)+\sin(6x) dx=-6\cos(x)-\frac{1}{6}\cos(6x)\]

OpenStudy (anonymous):

plug in \[\pi\] get \[-6\cos(\pi)-\frac{1}{6}\cos(6\pi)=6-\frac{1}{6}\]plug in 0 get \[-6\cos(0)-\frac{1}{6}\cos(0)=-6+\frac{1}{6}\] subtract the second from the first and get 12